# Monadic Second Order Logic of the Real Order with the Set Quantifier ranging over the Borel Sets

### v3

@@ 4.5. Subsets of $T^{1+\alpha}$ and branches in $T^{1+\alpha}]$
In [19, §2.3], Rabin showed how to consider MSO-formulas over the real order where only closed and $F_{\sigma}$-subsets of $2^{\omega}$ are allowed for the set variables, as being MSO-formulas over $T$. This translation, along with the decidability of S2S that he proved in his Main Theorem [19, Th 1.1] allowed him to show in [19, Th 2.17] that the MSO theory of the real order where only closed and $F_{\sigma}$-subsets of $2^{\omega}$ are allowed for the set variables, is decidable. Our main Theorem 5.1 states that the MSO theory of the real order where Borel sets are allowed for the set variables, is decidable. To prove it, we follow the reasoning of Rabin’s proof but we use subsets of $T^{1+\alpha}$ to represent $\Sigma_{\alpha}\cup\Pi_{\alpha}$-sets through the coding we developed. The most important remaining part of the translation is done through the Main Lemma 4.6. Before embarking to it, we develop a technique which will let us go back and forth between subsets of $T^{1+\alpha}$ and sets of branches in $T^{1+\alpha}$ in an MSO-definable way. This is the same technique that Rabin used for $T$ in [19, §2.3], only we apply it to $T^{1+\alpha}$. Let us develop it now.
@@ 5. Decidability of MSO over the real order with Borel Set Quantifiers
In the sentence above the statement of Conjecture 7A, Shelah motivates the conjectures by mentioning Rabin’s [19, Th 2.7], which he cites as “quantification over $F_{\sigma}$ sets gives a decidable theory”. The precise statement of [19, Th 2.7] assumes that the language has separate symbols for set variables which are closed sets and those that are $F_{\sigma}$-sets. The analogous language for Borel sets can be understood in two different ways, which turn out to have a different meaning. We shall first explain how to handle the weaker meaning, which is to work with the language that has separate symbols for set variables on each level of the Borel hierarchy. This can also be implemented by adding unary relation symbols $\Sigma_{\alpha}$ and $\Pi_{\alpha}$ acting on the set variables. We call the resulting interpretation of the Conjectures 7A/7B from [20] the Weak Shelah Conjecture. We confirm that conjecture in the following Theorem 5.1. We confirm the stronger version of the conjecture in Corollary 5.3. That stronger version is explained after the proof of Theorem 5.1.
@@
Theorem 5.1 does not answer the Strong Shelah’s Conjecture as the set of MSO-sentences possible to express in $\mathfrak{C}^{\ast}$ is strictly richer than the analogously defined set using $\mathfrak{C}$. An example of an MSO-sentence which is expressible in $\mathfrak{C}^{\ast}$ and not in $\mathfrak{C}$ is the statement of the Borel Determinacy (see §6). All known proofs of the Borel Determinacy involve an induction over the $\omega_{1}$-many levels of the Borel hierarchy and then making the final conclusion using the fact that if $\varphi(\alpha)$ is true for every $\alpha<\omega_{1}$ then $(\forall\alpha<\omega_{1})\varphi(\alpha)$ is true. This is not expressible in S2S. However, we can approach the Strong Shelah’s Conjecture using the fact that the MSO theory of $(\omega_{1},<)$ is decidable, as proven by Büchi in [3], and in a different way by Shelah in [20, Conclusion 4.10]. A consequence of this result is the following:
Theorem 5.2 (Büchi).
@@ 6. Effective Borel Determinacy
Clearly, a set is determined iff its complement is determined. It is quite easy to see that finite sets are determined. The following is the classical Gale-Stewart Theorem from [6], see [9, Th 20.1].
@@
Philip Wolfe [21] proved the determinacy of $F_{\sigma}$-sets. In [19, §2.5], Rabin showed that the statement that every $F_{\sigma}$-set is determined, is expressible as an MSO-sentence over ${\mathfrak{N}}_{2}$. He concluded that therefore, Wolfe’s theorem is among those decided by the decision procedure for S2S. In particular, knowing that all $F_{\sigma}$-sets are determined, we can conclude that for every $F_{\sigma}$-set there is a recursive procedure to decide which player wins the game. This recursive procedure is given by a run of the Rabin automaton. This is what we may call an effective version of the determinacy for $F_{\sigma}$-sets.
Borel Determinacy is a theorem of ZFC, see [9, Th 20.5], attributed to Donald Martin. Using the coding of arbitrary Borel sets we introduced, Rabin’s method of translating MSO-sentences to the runs of the Rabin automata and the fact that we know that Borel Determinacy is true, we obtain an effective Borel’s determinacy. Namely, for every Determinacy Game $G(H)$ where $H$ is a Borel set, there is a run of a Rabin automaton that determines the player that has the winning strategy. To summarise the situation:
@@
Proof.
On [19, pg.17] Rabin explains that strategies in the Determinacy Game can be viewed as characteristic functions of subsets $A$ of $T$. He deals with $F_{\sigma}$-sets, but for us, let us simply note that he gives MSO-formulas $W_{L}({\mathbf{A}},{\mathbf{D}})$ for $L\in\{I,II\}$ over $T$ which express that $\chi_{A}$ is a winning strategy for Player L in the Determinacy Game for an $F_{\sigma}$- or a $G_{\delta}$-set $D$. The formula $W_{L}({\mathbf{A}},{\mathbf{D}})$ is a conjunction of an MSO-formula $w_{L}({\mathbf{A}},{\mathbf{D}})$ and a separate formula stating that $\mathbf{D}$} is an $F_{\sigma}$-set or its complement.
@@ 7. Closing Remarks
We note that our use of the word “effective” in the statement of Theorem 6.3 is in the sense that the word is used in mathematical logic, namely that there is a theoretical computing machine which has a run that computes the problem in question. This does not say anything about an actual implementation of an algorithm for this computation. Such an implementation is clearly non-trivial, because Rabin automata work on trees with infinite branches, hence do not have a physical equivalent. Instead of having a physical Rabin automaton, one can construct a simulator, such as the Rabinizer [10], which, roughly speaking, to a given programme of the Rabin automaton associates an executable programme in a real programming language. (In fact Rabinizer can start from a formula of the linear temporal logic, translate it into a run of a deterministic Rabin automaton and execute it through a simulation in Java). In this context one can discuss what the computer scientists would understand by “effectiveness”, which is a polynomial or better bound on the complexity of the simulation programme in Java. We are not aware of any work studying such complexity even for the determinacy for $F_{\sigma}$-sets.
@@
The elegant proof of the Borel Determinacy given by Kechris in [9, Th, 20.5] relates to Martin’s paper [15]. Martin mentions in [15] that it was proposed by Yannis Moschovakis in his book [18] that such a proof should exist, although it was not found then. Before this proof, some earlier papers on the subject used a different technique. The idea was to associate to each Borel game $G$ a game $G^{\ast}$ on an open set in a different Polish space, so that if the player I or II has a winning strategy in $G^{\ast}$, then the player I or II respectively has a winning strategy in $G$. The Borel Determinacy then follows from the Gale-Stewart theorem on the determinacy of Determinacy Gales on open sets in Polish spaces [6] (this would require proper formulation of such games, more general than Definition 6.1). The idea of translation between $G$ and $G^{*}$ was introduced by Martin in [13], where he claims the proof of Borel Determinacy from the existence of an $\omega_{1}$-Erdős cardinal. His paper [14] claims Borel determinacy in ZFC, using a better coding. In the two papers the details of the proofs are missing or incorrect. With this in mind, we have tried to prove Borel Determinacy in yet another way, to match Martin’s original intuition. This would be by playing games on codes for $\Sigma_{\alpha}\cup\Pi_{\alpha}$-sets, understanding those as closed games in an appropriate Polish subspace of $T^{1+\alpha}]$ (see §2.1) and then projecting to the ordinary games on closed sets. The intended proof is by induction on $\alpha$ and uses a direct proof for $F_{\sigma}$-determinacy given in [9, Exercise 20.9] at the non-trivial part of the projection between games in the successor case. This proof would correspond to the original programme of proving Borel Determinacy as intended by Martin in his papers [13] and [14]. We have not worked out all of the details of this proof, as from the purely mathematical point of view it seems unnecessary to give a very long combinatorial argument for a theorem whose proof is already known and well written ([9, Th 20.5]). In fact, we believe that the combinatorial proof is simply a recounting of the decision procedure for the player which has the winning strategy. The part of the combinatorial proof that we did write is available in the code of this paper on the ArXiv. If an interested person uses it, they should ask for a permission and their work should have a proper mention of the source (this paper). The reason we mention these issues is that we believe that this algorithm might be of interest in the context of implementing it on the simulators of Rabin automata, such as Rabinizer [10].
@@
We finally mention that the proof of Borel Determinacy from [9, Th. 20.5] or [15] has indeed been implemented in Lean by Manthe [12]. Manthe has also published on ArXiv an outline of an intended proof of Theorem 5.1 [11] and has presented the same in talks, since October 2024. We would be eager to see the details of the proof if it converges, especially since its announcement uses a very different technique than the one presented here.

### v4

@@ 4.5. Subsets of $T^{1+\alpha}$ and branches in $T^{1+\alpha}]$
In [20, §2.3], Rabin showed how to consider MSO-formulas over the real order where only closed and $F_{\sigma}$-subsets of $2^{\omega}$ are allowed for the set variables, as being MSO-formulas over $T$. This translation, along with the decidability of S2S that he proved in his Main Theorem [20, Th 1.1] allowed him to show in [20, Th 2.17] that the MSO theory of the real order where only closed and $F_{\sigma}$-subsets of $2^{\omega}$ are allowed for the set variables, is decidable. Our main Theorem 5.1 states that the MSO theory of the real order where Borel sets are allowed for the set variables, is decidable. To prove it, we follow the reasoning of Rabin’s proof but we use subsets of $T^{1+\alpha}$ to represent $\Sigma_{\alpha}\cup\Pi_{\alpha}$-sets through the coding we developed. The most important remaining part of the translation is done through the Main Lemma 4.6. Before embarking to it, we develop a technique which will let us go back and forth between subsets of $T^{1+\alpha}$ and sets of branches in $T^{1+\alpha}$ in an MSO-definable way. This is the same technique that Rabin used for $T$ in [20, §2.3], only we apply it to $T^{1+\alpha}$. Let us develop it now.
@@ 5. Decidability of MSO over the real order with Borel Set Quantifiers
In the sentence above the statement of Conjecture 7A, Shelah motivates the conjectures by mentioning Rabin’s [20, Th 2.7], which he cites as ‘‘quantification over $F_{\sigma}$ sets gives a decidable theory’’. The precise statement of [20, Th 2.7] assumes that the language has separate symbols for set variables which are closed sets and those that are $F_{\sigma}$-sets. The analogous language for Borel sets can be understood in two different ways, which turn out to have a different meaning. We shall first explain how to handle the weaker meaning, which is to work with the language that has separate symbols for set variables on each level of the Borel hierarchy. This can also be implemented by adding unary relation symbols $\Sigma_{\alpha}$ and $\Pi_{\alpha}$ acting on the set variables. We call the resulting interpretation of the Conjectures 7A/7B from [21] the Weak Shelah Conjecture. This conjecture is also listed as an open question in [8, pg. 503]. We confirm the Weak Shelah’s conjecture in the following Theorem 5.1. We confirm the stronger version of the conjecture in Corollary 5.3. That stronger version is explained after the proof of Theorem 5.1.
@@
Theorem 5.1 does not answer the Strong Shelah’s Conjecture as the set of MSO-sentences possible to express in $\mathfrak{C}^{\ast}$ is strictly richer than the analogously defined set using $\mathfrak{C}$. An example of an MSO-sentence which is expressible in $\mathfrak{C}^{\ast}$ and not in $\mathfrak{C}$ is the statement of the Borel Determinacy (see §6). All known proofs of the Borel Determinacy involve an induction over the $\omega_{1}$-many levels of the Borel hierarchy and then making the final conclusion using the fact that if $\varphi(\alpha)$ is true for every $\alpha<\omega_{1}$ then $(\forall\alpha<\omega_{1})\varphi(\alpha)$ is true. This is not expressible in S2S. However, we can approach the Strong Shelah’s Conjecture using the fact that the MSO theory of $(\omega_{1},<)$ is decidable, as proven by Bьchi in [3], and in a different way by Shelah in [21, Conclusion 4.10]. A consequence of this result is the following:
Theorem 5.2 (Bьchi).
@@ 6. Effective Borel Determinacy
Clearly, a set is determined iff its complement is determined. It is quite easy to see that finite sets are determined. The following is the classical Gale-Stewart Theorem from [7], see [10, Th 20.1].
@@
Philip Wolfe [23] proved the determinacy of $F_{\sigma}$-sets. In [20, §2.5], Rabin showed that the statement that every $F_{\sigma}$-set is determined, is expressible as an MSO-sentence over ${\mathfrak{N}}_{2}$. He concluded that therefore, Wolfe’s theorem is among those decided by the decision procedure for S2S. In particular, knowing that all $F_{\sigma}$-sets are determined, we can conclude that for every $F_{\sigma}$-set there is a recursive procedure to decide which player wins the game. This recursive procedure is given by a run of the Rabin automaton. This is what we may call an effective version of the determinacy for $F_{\sigma}$-sets.
Borel Determinacy is a theorem of ZFC, see [10, Th 20.5], attributed to Donald Martin. Using the coding of arbitrary Borel sets we introduced, Rabin’s method of translating MSO-sentences to the runs of the Rabin automata and the fact that we know that Borel Determinacy is true, we obtain an effective Borel’s determinacy. Namely, for every Determinacy Game $G(H)$ where $H$ is a Borel set, there is a run of a Rabin automaton that determines the player that has the winning strategy. To summarise the situation:
@@
Доказательство.
On [20, pg.17] Rabin explains that strategies in the Determinacy Game can be viewed as characteristic functions of subsets $A$ of $T$. He deals with $F_{\sigma}$-sets, but for us, let us simply note that he gives MSO-formulas $W_{L}({\mathbf{A}},{\mathbf{D}})$ for $L\in\{I,II\}$ over $T$ which express that $\chi_{A}$ is a winning strategy for Player L in the Determinacy Game for an $F_{\sigma}$- or a $G_{\delta}$-set $D$. The formula $W_{L}({\mathbf{A}},{\mathbf{D}})$ is a conjunction of an MSO-formula $w_{L}({\mathbf{A}},{\mathbf{D}})$ and a separate formula stating that $\mathbf{D}$} is an $F_{\sigma}$-set or its complement.
@@ 7. Closing Remarks
We note that our use of the word ‘‘effective’’ in the statement of Theorem 6.3 is in the sense that the word is used in mathematical logic, namely that there is a theoretical computing machine which has a run that computes the problem in question. This does not say anything about an actual implementation of an algorithm for this computation. Such an implementation is clearly non-trivial, because Rabin automata work on trees with infinite branches, hence do not have a physical equivalent. Instead of having a physical Rabin automaton, one can construct a simulator, such as the Rabinizer [11], which, roughly speaking, to a given programme of the Rabin automaton associates an executable programme in a real programming language. (In fact Rabinizer can start from a formula of the linear temporal logic, translate it into a run of a deterministic Rabin automaton and execute it through a simulation in Java). In this context one can discuss what the computer scientists would understand by ‘‘effectiveness’’, which is a polynomial or better bound on the complexity of the simulation programme in Java. We are not aware of any work studying such complexity even for the determinacy for $F_{\sigma}$-sets.
@@
The elegant proof of the Borel Determinacy given by Kechris in [10, Th, 20.5] relates to Martin’s paper [16]. Martin mentions in [16] that it was proposed by Yannis Moschovakis in his book [19] that such a proof should exist, although it was not found then. Before this proof, some earlier papers on the subject used a different technique. The idea was to associate to each Borel game $G$ a game $G^{\ast}$ on an open set in a different Polish space, so that if the player I or II has a winning strategy in $G^{\ast}$, then the player I or II respectively has a winning strategy in $G$. The Borel Determinacy then follows from the Gale-Stewart theorem on the determinacy of Determinacy Gales on open sets in Polish spaces [7] (this would require proper formulation of such games, more general than Definition 6.1). The idea of translation between $G$ and $G^{*}$ was introduced by Martin in [14], where he claims the proof of Borel Determinacy from the existence of an $\omega_{1}$-Erdős cardinal. His paper [15] claims Borel determinacy in ZFC, using a better coding. In the two papers the details of the proofs are missing or incorrect. With this in mind, we have tried to prove Borel Determinacy in yet another way, to match Martin’s original intuition. This would be by playing games on codes for $\Sigma_{\alpha}\cup\Pi_{\alpha}$-sets, understanding those as closed games in an appropriate Polish subspace of $T^{1+\alpha}]$ (see §2.1) and then projecting to the ordinary games on closed sets. The intended proof is by induction on $\alpha$ and uses a direct proof for $F_{\sigma}$-determinacy given in [10, Exercise 20.9] at the non-trivial part of the projection between games in the successor case. This proof would correspond to the original programme of proving Borel Determinacy as intended by Martin in his papers [14] and [15]. We have not worked out all of the details of this proof, as from the purely mathematical point of view it seems unnecessary to give a very long combinatorial argument for a theorem whose proof is already known and well written ([10, Th 20.5]). In fact, we believe that the combinatorial proof is simply a recounting of the decision procedure for the player which has the winning strategy. The part of the combinatorial proof that we did write is available in the code of this paper on the ArXiv. If an interested person uses it, they should ask for a permission and their work should have a proper mention of the source (this paper). The reason we mention these issues is that we believe that this algorithm might be of interest in the context of implementing it on the simulators of Rabin automata, such as Rabinizer [11].
@@
We finally mention that the proof of Borel Determinacy from [10, Th. 20.5] or [16] has indeed been implemented in Lean by Manthe [13]. Manthe has also published on ArXiv an outline of an intended proof of Theorem 5.1 [12] and has presented the same in talks, since October 2024. We would be eager to see the details of the proof if it converges, especially since its announcement uses a very different technique than the one presented here.
Список литературы
  • [1] J. Richard Büchi. Weak second-order arithmetic and finite automata. Zeitschrift fur Mathematische Logik und Grundlagen der Mathematik, 6:66–92, 1960.
  • [2] J. Richard Büchi. Decision methods in the theory of ordinals. Bull. Amer. Math. Soc., 71:767–770, 1965.
  • [3] J. Richard Büchi. Decidable Theories II: The Monadic Second Order Theory of All Countable Ordinals, chapter The monadic second order theory of $\omega_{1}$, pages 1–127. Springer Berlin Heidelberg, Berlin, Heidelberg, 1973.
  • [4] Alonzo Church. An unsolvable problem of elementary number theory. American Journal of Mathematics, 58(2):345–363, 1936. URL: http://www.jstor.org/stable/2371045.
  • [5] Mirna Džamonja. A note on the splitting property in strongly dense posets of size $\aleph_{0}$. Rad. Mat., 8(2):321–326, 1998.
  • [6] Ryszard Engelking. General Topology. Heldermann, Berlin, 1989 (rev. ed.).
  • [7] David Gale and F. M. Stewart. Infinite games with perfect information. Contrib. Theory of Games, II, Ann. Math. Stud. No. 28, 245-266 (1953)., 1953.
  • [8] Yuri Gurevich. Monadic second-order theories. In Jon Barwise and Solomon Feferman, editors, Model-Theoretic Logics, pages 479–506. Springer-Verlag, 1985.
  • [9] Yuri Gurevich and Saharon Shelah. Monadic theory of order and topology in ZFC. Annals of Mathematical Logic, 23(2):179–198, 1982. URL: https://www.sciencedirect.com/science/article/pii/0003484382900043, doi:10.1016/0003-4843(82)90004-3.
  • [10] Alexander S. Kechris. Classical descriptive set theory, volume 156 of Graduate Texts in Mathematics. Springer-Verlag, New York, 1995.
  • [11] Jan Křetínský, Tobias Meggendorfer, Salomon Sickert, and Christopher Ziegler. Rabinizer 4: From LTL to your favourite deterministic automaton. In Hana Chockler and Georg Weissenbacher, editors, Computer Aided Verification, pages 567–577, Cham, 2018. Springer International Publishing.
  • [12] Sven Manthe. The Borel monadic theory of order is decidable, October 2024. URL: arXiv:2410.00887.
  • [13] Sven Manthe. A formalization of Borel determinacy in Lean, October (February for v1) 2025. URL: arXiv:2502.03432v2.
  • [14] Donald A. Martin. Measurable cardinals and analytic games. Fundam. Math., 66:287–291, 1970. URL: https://eudml.org/doc/214151, doi:10.4064/fm-66-3-287-291.
  • [15] Donald A. Martin. Borel determinacy. Ann. of Math. (2), 102(2):363–371, 1975. doi:10.2307/1971035.
  • [16] Donald A. Martin. A purely inductive proof of Borel determinacy. Recursion theory, Proc. AMS-ASL Summer Inst., Ithaca/N.Y. 1982, Proc. Symp. Pure Math. 42, 303-308 (1985)., 1985.
  • [17] Yu. V. Matiyasevich. Enumerable sets are diophantine. Sov. Math., Dokl., 11:354–358, 1970.
  • [18] Robert McNaughton. Testing and generating infinite sequences by a finite automaton. Inf. Control, 9:521–530, 1966. doi:10.1016/S0019-9958(66)80013-X.
  • [19] Yiannis N. Moschovakis. Descriptive set theory, volume 100 of Stud. Logic Found. Math. Elsevier, Amsterdam, 1980.
  • [20] Michael O. Rabin. Decidability of second order theories and automata on infinite trees. Trans. Amer. Math. Soc., 149:1–35, 1969.
  • [21] Saharon Shelah. The monadic theory of order. Ann. of Math. (2), 102(3):379–419, 1975. doi:10.2307/1971037.
  • [22] Boris A. Trakhtenbrot. The synthesis of logical nets whose operators are described in terms of monadic predicates. Doklady Akademii Nauk SSSR, 118(4):646–649, 1958. In Russian.
  • [23] Philip Wolfe. The strict determinateness of certain infinite games. Pac. J. Math., 5:841–847, 1955. doi:10.2140/pjm.1955.5.841.